OpenStudy (anonymous):

what is the answer for 3/4+(negative)1/2

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

\[\frac{3}{4}-\frac{1}{2}\] \[=\frac{3-2}{4}\] \[=\frac{1}{4}\]